OpenStudy (anonymous):

i have a question on my math, not sure what to do to get rid of the fractions. Do i work in reverse? Help! There was a jar of cookies. Amanda ate 1/2 of the cookies. Beth came and ate 1/3 of the cookies. Christine came and ate 1/4 of the cookies. Daniel came and ate 1 cookie. When Eva looked in the jar there were only 2 cookies left. How many cookies were there to begin with?

OpenStudy (lgbasallote):

what do you mean reverse exactly?

OpenStudy (anonymous):

working this problem out in reverse.

OpenStudy (lgbasallote):

well what you do is just multiply the fractions

OpenStudy (lgbasallote):

\[ x \times \frac12 \times \frac 13 \times \frac 14 - 1 = 2\] where x was the original number of cookies does that help?

OpenStudy (anonymous):

what number do I multiply the fractions? this is where I'm getting confused

OpenStudy (lgbasallote):

what do you mean what number? you just multiply the fractions

OpenStudy (lgbasallote):

for example \[\frac 12 \times \frac 12 \implies \frac{1\times 1}{2\times 2} \implies \frac 14\]

OpenStudy (anonymous):

so when I multiplied straight through i got 1/24x=2, is that correct or did I do something wrong

OpenStudy (anonymous):

so x=12

OpenStudy (lgbasallote):

well close but no...you should get \[\frac{x}{24}-1 = 2\] so now you add 1 to both sides \[\frac{x}{24} = 3\] then multiply both sides by 24 to get x do you get it?

OpenStudy (anonymous):

oh I forgot about the minus 1. I just worked it out and got 72, correct?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

THANK YOU! I understand it now

OpenStudy (lgbasallote):

welcome